OpenStudy (anonymous):

a line contains the point (8,-5). if the slope of the line is 5/7, write the equation of the line using point slope form. a. y-5=5/7 (x+8) b. y + 5 = -5/8 (x+7) c. y+8= 5/7( x-5) d. y+5= 5/7(x-8)

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

Member what we did in the previous post? It is the same thing except that the slope is not 0, it is 5/7. And the point is already give, which is (8,-5) So, this equation is perfect for two methods, the point slope formula or slope intercept form. For point slope formula: \(y-\color{red}{y_1 }=m(x-\color{red}{x_1})\) \(y+5=\frac{5}{7}(x-8)\) For slope intercept form: \(y=mx+b\) Plug everything in x,y=8,-5 m=5/7 and solve for b \(-5=\frac{5}{7}(5)+b\) and then rewrite the equation with m and slope numbers plugged in.

OpenStudy (zale101):

\(-5=\frac{5}{7}(8)+b\) The previous one was an error, sorry for that.

OpenStudy (anonymous):

so how do i solve it?? do i multiply??

OpenStudy (zale101):

But everything was correct

OpenStudy (anonymous):

oh so its a right?!?!?!

OpenStudy (zale101):

You solve the equation from which one looks easier to you. I find this easier for algebra students, it's pretty straight forward \(y-(-5)=\frac{5}{7}(x-8)\)

OpenStudy (anonymous):

so its not a??

OpenStudy (zale101):

It can't be a, it is not the same as this \(y+5=\frac{5}{7}(x-8)\)

OpenStudy (zale101):

And lol, i didnt pay attention to the choices, i would've just choose the first method by using the point slope formula

OpenStudy (zale101):

So which is the right choice?

OpenStudy (anonymous):

a?

OpenStudy (zale101):

This choice ? ---> a. y-5=5/7 (x+8)

OpenStudy (anonymous):

no it was incorrect. have no idea then

OpenStudy (anonymous):

oh ok d then (;

OpenStudy (zale101):

Yes

OpenStudy (anonymous):

thx

OpenStudy (zale101):

np, but you already understand how to do these kind of problems now, right?

OpenStudy (anonymous):

yes i suppose